during the first part of a trip a canoeist travels 24 miles at a certain speed. The canoeist travels 7 miles on the second part of the trip at a speed 5 miles per hour slower. The total time for the trip is 3 hours . What was the speed for each part of the trip?
Answer by TimothyLamb(4379)

x = speed first part
y = speed second part = x - 5
---
s = d/t
t = d/s
---
first part:
a = 24/x
---
second part:
t = 7/y
b = 7/(x - 5)
---
a + b = 3
24/x + 7/(x - 5) = 3
24(x - 5)/x(x - 5) + 7x/x(x - 5) = 3
24(x - 5) + 7x = 3x(x - 5)
24x - 5*24 + 7x = 3xx - 15x
24x - 120 + 7x = 3xx - 15x
3xx - 15x - 24x - 7x + 120 = 0
3xx - 46x + 120 = 0
---
the above quadratic equation is in standard form, with a=3, b=-46 and c=120
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
3 -46 120
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 12
x = 3.33333333
---
the root x=3.333 doesn't fit the problem statement, because (y=3.333 - 5) and y becomes negative, so use the root x=12:
---
x = 12
y = x - 5
y = 12 - 5
y = 7
---
answer:
x = speed first part = 12 mph
y = speed second part = 7 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php